OpenStudy (anonymous):

mike must build a rectangular storage shed, he wants the length to be 6ft greater than the width, and the perimeter will be 44ft. find the length and with of the shed?

OpenStudy (mathstudent55):

Let L = length Let W = width he wants the length to be 6ft greater than the width: L = W + 6 Perimeter = P = L + L + W + W = 2L + 2W 2L + 2W = 44, but L = W + 6, so 2(W + 6) + 2W = 44 2W + 12 + 2W = 44 4W + 12 = 44 4W = 32 W = 8 L = W + 6 = 14